# CaberNet: Causal Representation Learning for Cross-Domain HVAC Energy Prediction

### v1

@@ Abstract.
Cross-domain HVAC energy prediction is essential for scalable building energy management, particularly because collecting extensive labeled data for every new building is both costly and impractical. Yet, this task remains highly challenging due to the scarcity and heterogeneity of data across different buildings, climate zones, and seasonal patterns. In particular, buildings situated in distinct climatic regions introduce variability that often leads existing methods to overfit to spurious correlations, rely heavily on expert intervention, or compromise on data diversity. To address these limitations, we propose CaberNet, a causal and interpretable deep sequence model that learns invariant (Markov blanket) representations for robust cross-domain prediction. In a purely data-driven fashion and without requiring any prior knowledge, CaberNet integrates i) a global feature gate trained with a self-supervised Bernoulli regularization to distinguish superior causal features from inferior ones, and ii) a domain-wise training scheme that balances domain contributions, minimizes cross-domain loss variance, and promotes latent factor independence. We evaluate CaberNet on real-world datasets collected from three buildings located in three climatically diverse cities, and it consistently outperforms all baselines, achieving a 22.9% reduction in normalized mean squared error (NMSE) compared to the best benchmark. Our code is available at https://github.com/rickzky1001/CaberNet-CRL.
@@ 6. Experiments
Datasets. Publicly available datasets that combine indoor sensing with HVAC energy consumption across multiple climate zones remain scarce. Therefore, our experiments are conducted on data we collected via multi-site sensor deployments. The data comprises six office-floor datasets from three commercial buildings located in three Chinese cities, spanning a north–middle–south climate gradient (Figure 5). Indoor sensing is deployed uniformly across the occupied areas. Per timestamp, the feature set includes: indoor temperature, indoor humidity, light intensity, $\mathrm{CO_{2}}$ concentration, indoor air pressure, and total volatile organic compounds (TVOC). Indoor sensors log at a 5-minute resolution. We also ingest outdoor temperature collected from NASA POWER (nasa_power) and resample it to 5-minute resolution via linear interpolation to match the indoor data cadence. The data used in the experiments cover the period from November 2024 to July 2025.
@@ 6.1. Performance Comparison
Method Beijing Hefei Huizhou Average
L 22 L 23 L 9 L 10 L 15 L 16
@@
Shift-GRU achieves the second average NMSE and is especially strong in Huizhou, where it beats plain LSTM, likely because per-feature time-lag alignment mitigates phase shifts from thermal inertia and lag offsets in warm-humid climates. However, Pearson correlation is linear and pairwise, so its shifts can be unreliable under nonlinear or multi-lag dynamics. This limits robustness relative to CaberNet. By contrast, STL performs worst overall: its similarity-based filtering shrinks data diversity and sample size, which is ill-suited when data are scarce and highly variable across domains.
LSTM–LiNGAM is overall comparable to LSTM, and in some cases markedly better (Beijing L22 and Huizhou L16), probably because the discovered blanket retained strong drivers (e.g., outdoor_temperature, is_work) while discarding weak ones, as shown in Appendix B. However, its performance varies: hard masking can remove useful proxy variables; LiNGAM also assumes a linear, acyclic data-generating process with non-Gaussian errors and no latent confounding (shimizu2011directlingam), which may be too restrictive for the complex and nonlinear nature of HVAC dynamics. Moreover, LiNGAM’s binary include/exclude decision is brittle: features that are not strictly causal but still carry a latent causal signal (e.g., proxies, mediators, regime indicators) are dropped entirely. By contrast, CaberNet uses soft gating to retain such partial contributions while appropriately down-weighting them. In this way, the model can still exploit these signals in forming the final causal representation, thereby improving robustness under domain shift. These misspecifications can yield suboptimal blankets and underperformance on other floors. Notably, in Hefei L10 LiNGAM selected essentially the full feature set as the blanket, making its input identical to LSTM and resulting in identical NMSE. The detailed SCMs constructed by LiNGAM are shown in Appendix B.
@@ 6.2.1. Regularization
Method Beijing Hefei Huizhou Average
L 22 L 23 L 9 L 10 L 15 L 16
@@
LSTM (ERM) vs. CaberNet-SIRM (IRM). Compared to the LSTM baseline (Table 1), ERM can occasionally win on a specific test domain (e.g., Beijing L23). This happens because ERM’s sample-proportional training implicitly upweights large/source-like domains; when such a domain is “lucky” and happens to resemble the held-out target distribution, ERM benefits from this alignment. By contrast, CaberNet-SIRM gives each training domain equal influence, foregoing that accidental advantage. This “ERM>>IRM” phenomenon is common, especially when ERM is well tuned (rosenfeld2020risks; gulrajani2020search); nonetheless, on average CaberNet-SIRM still outperforms ERM by 12.7%, indicating that equalizing domain contributions reduces overfitting to domain-specific variations of large domains and improves robustness across diverse targets.
@@ Appendix B Causal Graphs Extracted by LSTM-LiNGAM
BJ22 {IC, OT} {W, IP, IL, IH} {IL, IC, IH, W, IP, OT}
BJ23 {IT, IL, IC, IP, OT, W} {IT, IL, IC, OT, W, IP}
HF9 {OT} {IP, IL} {IP, OT, IL}
HF10 {IV, IP, W} {IT, IH, IC, OT} {W, IV, OT, IP, IT, IL} {W, IV, IC, OT, IP, IL, IT, IH}
HZ15 {IT, IL, IP} {IH, IV} {IV, IL, W, IP, IC, IT} {IH, IL, W, IT, IP, IC, IV}
HZ16 {IL, IP} {IH, OT} {IV, IT, IP, OT, IL} {IV, IT, IH, IP, OT, IL}
@@ Appendix C Detail of indy loss ablation study
Beijing L22 0.145 0.081 0.238 0.129 0.205 0.143 0.207 0.192
L23 0.149 0.142 0.162 0.162 0.161 0.159
Hefei L9 0.135 0.138 0.160 0.147 0.161 0.135 0.232 0.210
L10 0.263 0.238 0.278 0.263 0.323 0.307
Huizhou L15 0.299 0.285 0.279 0.279 0.280 0.261 0.237 0.214
L16 0.251 0.270 0.277 0.282 0.289 0.278

### v3

@@ Abstract.
Cross-domain HVAC energy prediction is essential for scalable building energy management, particularly because collecting extensive labeled data for every new building is both costly and impractical. Yet, this task remains highly challenging due to the scarcity and heterogeneity of data across different buildings, climate zones, and seasonal patterns. In particular, buildings situated in distinct climatic regions introduce variability that often leads existing methods to overfit to spurious correlations, rely heavily on expert intervention, or compromise on data diversity. To address these limitations, we propose CaberNet, a causal and interpretable deep sequence model that learns invariant (Markov blanket) representations for robust cross-domain prediction. In a purely data-driven fashion and without requiring any prior knowledge, CaberNet integrates i) a global feature gate trained with a self-supervised Bernoulli regularization to distinguish superior causal features from inferior ones, and ii) a domain-wise training scheme that balances domain contributions, minimizes cross-domain loss variance, and promotes latent factor independence. We evaluate CaberNet on real-world datasets collected from three buildings located in three climatically diverse cities, and it consistently outperforms all baselines, achieving a 22.9% reduction in normalized mean squared error (NMSE) compared to the best benchmark. Our code is available at https://github.com/SusCom-Lab/CaberNet-CRL.
@@ 6. Experiments
Datasets. Publicly available datasets that combine indoor sensing with HVAC energy consumption across multiple climate zones remain scarce. Therefore, our experiments are conducted on data we collected via multi-site sensor deployments. The data comprises six office-floor datasets from three commercial buildings located in three Chinese cities, spanning a north–Central–south climate gradient (Figure 5). Indoor sensing is deployed uniformly across the occupied areas. Per timestamp, the feature set includes: indoor temperature, indoor humidity, light intensity, $\mathrm{CO_{2}}$ concentration, indoor air pressure, and total volatile organic compounds (TVOC). Indoor sensors log at a 5-minute resolution. We also ingest outdoor temperature collected from NASA POWER (nasa_power) and resample it to 5-minute resolution via linear interpolation to match the indoor data cadence. The data used in the experiments cover the period from November 2024 to July 2025.
@@ 6.1. Performance Comparison
Method North Central South Average
L-A L-B L-A L-B L-A L-B
@@
Shift-GRU achieves the second average NMSE and is especially strong in the South, where it beats plain LSTM, likely because per-feature time-lag alignment mitigates phase shifts from thermal inertia and lag offsets in warm-humid climates. However, Pearson correlation is linear and pairwise, so its shifts can be unreliable under nonlinear or multi-lag dynamics. This limits robustness relative to CaberNet. By contrast, STL performs worst overall: its similarity-based filtering shrinks data diversity and sample size, which is ill-suited when data are scarce and highly variable across domains.
LSTM–LiNGAM is overall comparable to LSTM, and in some cases markedly better (North L-A and South L-B), probably because the discovered blanket retained strong drivers (e.g., outdoor_temperature, is_work) while discarding weak ones, as shown in Appendix B. However, its performance varies: hard masking can remove useful proxy variables; LiNGAM also assumes a linear, acyclic data-generating process with non-Gaussian errors and no latent confounding (shimizu2011directlingam), which may be too restrictive for the complex and nonlinear nature of HVAC dynamics. Moreover, LiNGAM’s binary include/exclude decision is brittle: features that are not strictly causal but still carry a latent causal signal (e.g., proxies, mediators, regime indicators) are dropped entirely. By contrast, CaberNet uses soft gating to retain such partial contributions while appropriately down-weighting them. In this way, the model can still exploit these signals in forming the final causal representation, thereby improving robustness under domain shift. These misspecifications can yield suboptimal blankets and underperformance on other floors. Notably, in Central L-B LiNGAM selected essentially the full feature set as the blanket, making its input identical to LSTM and resulting in identical NMSE. The detailed SCMs constructed by LiNGAM are shown in Appendix B.
@@ 6.2.1. Regularization
Method North Central South Average
L-A L-B L-A L-B L-A L-B
@@
LSTM (ERM) vs. CaberNet-SIRM (IRM). Compared to the LSTM baseline (Table 1), ERM can occasionally win on a specific test domain (e.g., North L-B). This happens because ERM’s sample-proportional training implicitly upweights large/source-like domains; when such a domain is “lucky” and happens to resemble the held-out target distribution, ERM benefits from this alignment. By contrast, CaberNet-SIRM gives each training domain equal influence, foregoing that accidental advantage. This “ERM>>IRM” phenomenon is common, especially when ERM is well tuned (rosenfeld2020risks; gulrajani2020search); nonetheless, on average CaberNet-SIRM still outperforms ERM by 12.7%, indicating that equalizing domain contributions reduces overfitting to domain-specific variations of large domains and improves robustness across diverse targets.
@@ Appendix B Causal Graphs Extracted by LSTM-LiNGAM
North L-A {IC, OT} {W, IP, IL, IH} {IL, IC, IH, W, IP, OT}
North L-B {IT, IL, IC, IP, OT, W} {IT, IL, IC, OT, W, IP}
Central L-A {OT} {IP, IL} {IP, OT, IL}
Central L-B {IV, IP, W} {IT, IH, IC, OT} {W, IV, OT, IP, IT, IL} {W, IV, IC, OT, IP, IL, IT, IH}
South L-A {IT, IL, IP} {IH, IV} {IV, IL, W, IP, IC, IT} {IH, IL, W, IT, IP, IC, IV}
South L-B {IL, IP} {IH, OT} {IV, IT, IP, OT, IL} {IV, IT, IH, IP, OT, IL}
@@ Appendix C Detail of indy loss ablation study
North L-A 0.145 0.081 0.238 0.129 0.205 0.143 0.207 0.192
L-B 0.149 0.142 0.162 0.162 0.161 0.159
Central L-A 0.135 0.138 0.160 0.147 0.161 0.135 0.232 0.210
L-B 0.263 0.238 0.278 0.263 0.323 0.307
South L-A 0.299 0.285 0.279 0.279 0.280 0.261 0.237 0.214
L-B 0.251 0.270 0.277 0.282 0.289 0.278